OpenStudy (anonymous):

which is the vertex of the graph of y=x^2+5x+6

OpenStudy (jdoe0001):

\(\bf y=x^2+5x+6\implies y = {\color{red}{ 1}}x^2{\color{blue}{+5}}x{\color{green}{ +6}}\\ \textit{vertex of a parabola}\\ \quad \\ \left(-\cfrac{{\color{blue}{ b}}}{2{\color{red}{ a}}}\quad ,\quad {\color{green}{ c}}-\cfrac{{\color{blue}{ b}}^2}{4{\color{red}{ a}}}\right)\)

OpenStudy (anonymous):

Thank you!